# SiLVR: Scalable Lidar-Visual Radiance Field Reconstruction with Uncertainty Quantification

### v1

@@ V-B Implementation Details
When collecting the data, we use VILENS [16], a lidar-inertial SLAM system running online to estimate a globally consistent trajectory and to motion correct the lidar measurements. The SLAM trajectory estimated online can also be further optimised using Bundle Adjustment as described in Sec. IV-D1. This improves the visual reconstruction quality, as shown later in Tab. IV. Individual lidar scans are projected to form a sparse depth image coinciding with the camera image (i.e. the same camera pose and intrinsic parameters). Lidar point normals are also estimated using the lidar range image, and projected as a sparse normal image. We use the calibrations provided by the Oxford Spires dataset [19]. In this dataset, the intrinsics and extrinsics of the set of cameras are estimated using Kalibr [60], and a single extrinsic transformation between the three cameras and the lidar is estimated using DiffCal [61]. When running COLMAP, we further optimise the camera intrinsics produced by Kalibr. To train the NeRF model, we used an Nvidia RTX 3080 Ti. One training iteration takes 4096 rays.
Figure 7: Comparison of reconstruction quality of our method and other baseline methods in two real-world large-scale scenes. Among the radiance field baseline methods, SiLVR’s reconstruction is the most accurate and complete, especially on the ground where there is insufficient visual constraints.
@@ V-C2 Map Filtering for Fair Evaluation
Perfect accuracy and completeness scores (of zero in both cases) would be achieved if the two point clouds are identical, and any deviation is penalised by a higher value. In practice, the ground truth model and the reconstruction do not perfectly overlap, as they scan slightly different parts of the scene from different viewpoints. Two situations can occur which do not correspond to mapping error:
@@
    Missing regions in the ground truth map: the ground truth map can have undetected areas of the scene that were captured in the Frontier data sequence. This would lead to an artificially higher accuracy score for the Frontier data in such regions. These are in effect false positives.
@@
    Extra regions in the ground truth map: the ground truth map can contain areas that the Frontier device did not scan. In this case, the NeRF reconstruction of these extra regions will be missing. These are undesirable false negatives, and the completeness score would again be artificially higher than it should be.
@@
To address this issue, we filter the non-overlapping regions that we consider should not be included in the evaluation — for both the reconstruction and the ground truth. Specifically, our evaluation system first builds an occupancy map of the ground truth reconstruction using Octomap [28]. We then remove points in the reconstruction that are not in the octree (i.e. in the unknown space). Similarly, we build an occupancy map of the lidar point clouds, and remove ground truth points within the unknown space. Manual filtering is also applied for regions that are inside the buildings.
@@ VI-A Evaluation of the 3D Reconstruction
We summarise the quantitative results in Tab. II and Tab. III, and show the 3D reconstructions in Fig. 5 and Fig. 7. Among all methods, lidar-only method VILENS-SLAM is the most accurate and complete. OpenMVS produces much more accurate and complete reconstructions compared to the radiance field reconstruction, but produces a poor reconstruction of the ground compared to lidar-based methods. This is expected as there is little texture on the ground. All the radiance field reconstructions are less accurate and less complete compared to VILENS-SLAM and OpenMVS. Nerfacto fails to estimate most of the ground geometry accurately in ROQ (the reconstruction is below the ground and is filtered by the occupancy map described in Sec. V-C2). Compared to Nerfacto, NeuSfacto reconstructs the ground surface better, but still at an incorrect height compared to the ground truth. This shows that while the SDF formulation poses a geometric prior on the scene (enforcing that there should be a surface rather than an arbitrary volumetric field), it still cannot estimate the surface accurately if there are insufficient visual constraints. The 3D Gaussians exported by Splatfacto also cannot reconstruct the ground accurately. These 3D Gaussians are located mostly on the visual features of the sites—since they are initialised by the COLMAP feature points.
Compared to the vision-only methods, SiLVR incorporates lidar measurements and has significantly better reconstruction fidelity especially on the ground. Compared to VILENS-SLAM, SiLVR achieves more complete reconstruction for the shorter sequences (e.g., HBAC in Fig. 5) since it uses dense visual information. When there are many accumulated lidar points (which is the case for the large-scale datasets in Fig. 7), this advantage is less prominent.
@@ VI-B Evaluation of Epistemic Uncertainty Estimation
We evaluate the epistemic uncertainty estimates using both synthetic data and real-world data. The synthetic data is generated using the CARLA simulator [66], which provides perfect pose trajectories and ground truth maps. We simulated a vehicle with a lidar and three-cameras in a configuration similar to the Frontier. Meanwhile, the real-world dataset used in this section is the Radcliffe Observatory Quarter (ROQ).
@@ VI-B1 Evaluation Metrics
We evaluate the epistemic uncertainty estimates using the Sparsification Plot which has been used for evaluating confidence estimates in the literature [67, 68, 15]. The Sparsification Plot is used to evaluate how the uncertainty estimates coincide with the actual errors (in our case, point-to-point distance to the ground truth). In these plots, reconstruction points with the highest uncertainty are gradually removed, and the average errors of the remaining reconstruction points are calculated to form a graph. If the uncertainty estimates align perfectly with the actual errors, then the reconstruction points with the the highest errors are always removed first, which leads to the steepest decrease of the remaining error as the most uncertain points are being removed. This ideal sparsification curve is referred to as Oracle Sparsification. In practice, the uncertainty estimates do not align with the actual errors perfectly. Specifically, when a reconstruction point has a higher uncertainty but lower error compared to another point, the uncertainty estimates are considered not perfect. The area between the sparsification and its oracle indicates how different the uncertainty estimates are from the ideal ones, and can then be used to compute the Area Under Sparsification Error (AUSE). A smaller difference between the sparsification and its oracle results in a lower AUSE, and indicates that the uncertainty estimates are better because they align better with the actual errors.
@@ VI-B2 Results
Figure 9: Comparison between visual and lidar uncertainty. The reconstructions at the top are coloured by uncertainty estimates (red is high uncertainty, green is low uncertainty). Here, the lidar used was a Hesai XT32 which has a narrow field-of-view and cannot scan the upper part of the buildings, and the depth uncertainty estimates can identify such regions (red point clouds). This indicates that only visual information is used when reconstructing these areas.
The advantage of the disentangled uncertainty estimates is also demonstrated in Fig. 9 where we use a narrow vertical field-of-view lidar with wider vertical field-of-view cameras. Here, the lidar is not able to scan the upper part of the two buildings highlighted in 9, but the cameras can. Because of this, when we compute the epistemic uncertainty, we can see high lidar depth uncertainty for the upper part of both buildings. This indicates that the reconstruction is derived primarily from the visual data. In summary, from a sensor fusion point of view, our epistemic uncertainty estimation framework provides a systematic analysis of each sensor’s contribution to the final reconstruction.
Figure 10: Submapping strategy comparison. Visibility information guides the clustering algorithm to group images looking at the same object together, which then leads to more accurate and complete reconstruction of that object than algorithms that only consider distances.
@@ VI-C View Selection Strategy
In this section, we compare our visibility-based submapping strategy with an alternative distance-based submapping strategy proposed in [12]. As shown in Fig. 10, the building highlighted is divided into two submaps when using the distance-based submapping method. This is not ideal as it reduces the number of view constraints, which makes each submap’s partial reconstruction of that building have a lower quality. In addition, distance-based submapping put the poses in A and C into the same submap, which is in fact not ideal. While pose A and pose C are spatially close, they have opposite viewing directions: pose A is looking at the highlighted building, while pose C is looking away from it. In comparison, visibility-based submapping moves pose A into the submap that contains the highlighted building, and pose C into another submap.
@@ VI-D Multi-Camera Setup Ablation Study
The advantage of our multi-camera sensor setup is demonstrated qualitatively in Fig. 11. Compared to the three-camera setup, using only the front-facing camera leads to a reconstruction that is not only incomplete, but also with poorer geometry. Visual reconstruction with the photometric loss is limited to generating a good quality rendering only at the input viewing angle. The reconstruction using the front-only camera in Fig. 11 is trained with images looking in a single direction in the scene. This results in a poor geometric reconstruction when rendered from an unseen angle. In comparison, reconstruction with three cameras generates a more complete and more accurate reconstruction.
Figure 11: Comparison of reconstruction of HBAC building using the front camera only vs. using all the three cameras. The three-camera setup generates more complete and accurate reconstructions compared to using only a single front-facing camera. The multi-camera setting is important in robotic applications where it would be infeasible to actively scan the entire scene to obtain strong viewpoint constraints.
@@ VI-E Effect of Bootstrapping SLAM Poses
The results are summarised in Tab. IV. For all COLMAP configurations, providing the SLAM prior poses not only accelerates pose computation, but also leads to better test rendering, compared to COLMAP without any initialisation. Our SLAM prior poses can also register all the images in the trajectory; meanwhile COLMAP on its own only registers only 55%-95% images. Extracting more visual features per image (from 1024 to 8192) leads to a higher percentage of image registration and better visual reconstruction (PSNR and SSIM). This comes at the expense of a higher computation time, especially with the VocabTree matcher. Using the COLMAP Sequential Matcher is generally faster than Vocabulary Tree Matcher.

### v3

@@ V-B Implementation Details
V-B1 Data Collection and Processing
When collecting the data, we use VILENS [23], a lidar-inertial SLAM system running online to estimate a globally consistent trajectory and to motion correct the lidar measurements. The SLAM trajectory estimated online can also be further optimised using Bundle Adjustment as described in Sec. IV-D1. This improves the visual reconstruction quality, as shown later in Tab. IV. Individual lidar scans are projected to form a sparse depth image coinciding with the camera image (i.e., the same camera pose and intrinsic parameters). Surface normals of the lidar points are estimated as the surface normal of the polygon formed by the current point’s neighbouring points in the lidar range image using Newell’s method, and projected as a sparse normal image. We use the calibrations provided by the Oxford Spires dataset [26]. In this dataset, the intrinsics and extrinsics of the set of cameras are estimated using Kalibr [66], and a single extrinsic transformation between the three cameras and the lidar is estimated using DiffCal [67]. When running COLMAP, we further optimise the camera intrinsics produced by Kalibr.
V-B2 NeRF Reconstruction
Our NeRF reconstruction system extends Nerfacto, which is a specific vision-only pipeline implemented within the Nerfstudio framework [15]. Nerfacto’s rendering quality is comparable to state-of-the-art methods such as MipNeRF-360 [50] while achieving a substantial acceleration in reconstruction speed as it also incorporates efficient hash encoding which was proposed by the authors of Instant-NGP [12]. The scene contraction, proposed in [50], is also used to improve memory efficiency and to represent scenes with high-resolution content near the input camera locations. The contraction function non-linearly maps any point in space into a cube of side length 2, and represents the scene within this contracted space. In real-world outdoor environments, there are often large variations in exposure and lighting conditions. Because of this, we use a per-frame appearance encoding for each image, similar to [5, 51]. To train the NeRF model, we used an Nvidia RTX 4090. One training iteration takes 4096 rays.
Figure 6: Comparison of reconstruction quality of our method and other baseline methods in two real-world large-scale scenes. Among the radiance field baseline methods, SiLVR’s reconstruction is the most accurate and complete, especially on the ground where there are insufficient visual constraints.
@@ V-C2 Map Filtering for Fair Evaluation
Perfect accuracy and completeness scores (of zero in both cases) would be achieved if the two point clouds were identical, and any deviation is penalised by a higher value. In practice, the ground truth model and the reconstruction do not perfectly overlap, as they scan slightly different parts of the scene from different viewpoints. Two situations can occur which do not correspond to mapping error:
@@
    Missing regions in the ground truth map: the ground truth map can have undetected areas of the scene that were captured in the Frontier data sequence. This would lead to an artificially higher accuracy score for the Frontier data in such regions. These are in effect false positives as shown in Fig. 7.
@@
    Extra regions in the ground truth map: the ground truth map can contain areas that the Frontier device did not scan. In this case, the NeRF reconstruction of these extra regions will be missing. These are undesirable false negatives as shown in Fig. 7, and the completeness score would again be artificially higher than it should be.
@@
To address this issue, we filter the non-overlapping regions that we consider should not be included in the evaluation — for both the reconstruction and the ground truth. Specifically, our evaluation system first builds an occupancy map of the ground truth reconstruction using Octomap [35]. We then remove points in the reconstruction that are not in the octree (i.e., in the unknown space). Similarly, we build an occupancy map of the lidar point clouds, and remove ground truth points within the unknown space. Manual filtering is also applied for regions that are inside the buildings.
@@ VI-A Evaluation of the 3D Reconstruction
We summarise the quantitative results in Tab. II and Tab. III, and show the 3D reconstructions in Fig. 5 and Fig. 6. Among all methods, the lidar-only method VILENS-SLAM is the most accurate and complete. OpenMVS produces much more accurate and complete reconstructions compared to the radiance field reconstruction, but produces a poor reconstruction of the ground compared to lidar-based methods. This is expected as there is little texture on the ground. All the radiance field reconstructions are less accurate and less complete compared to VILENS-SLAM and OpenMVS. Nerfacto fails to estimate most of the ground geometry accurately in ROQ (the reconstruction is below the ground and is filtered by the occupancy map described in Sec. V-C2). Compared to Nerfacto, NeuSfacto reconstructs the ground surface better, but still at an incorrect height compared to the ground truth. This shows that while the SDF formulation poses a geometric prior on the scene (enforcing that there should be a surface rather than an arbitrary volumetric field), it still cannot estimate the surface accurately if there are insufficient visual constraints. The 3D Gaussians exported by Splatfacto also cannot reconstruct the ground accurately. These 3D Gaussians are located mostly on the visual features of the sites—since they are initialised by the COLMAP feature points.
Compared to the vision-only methods, SiLVR incorporates lidar measurements and has significantly better reconstruction fidelity, especially on the ground. Compared to VILENS-SLAM, SiLVR achieves more complete reconstruction for the shorter sequences (e.g., HBAC in Fig. 5) since it uses dense visual information. When there are many accumulated lidar points (which is the case for the large-scale datasets in Fig. 6), this advantage is less prominent.
@@ VI-B Evaluation of Epistemic Uncertainty Estimation
We evaluate the epistemic uncertainty estimates using both synthetic data and real-world data. The synthetic data is generated using the CARLA simulator [72], which provides perfect pose trajectories and ground truth maps. We simulated a vehicle with a lidar and three cameras in a configuration similar to the Frontier. Meanwhile, the real-world dataset used in this section is the Radcliffe Observatory Quarter (ROQ).
@@ VI-B1 Evaluation Metrics
We evaluate the epistemic uncertainty estimates using the Sparsification Plot which has been used for evaluating confidence estimates in the literature [73, 74, 22]. The Sparsification Plot is used to evaluate how the uncertainty estimates coincide with the actual errors (in our case, point-to-point distance to the ground truth). In these plots, reconstruction points with the highest uncertainty are gradually removed, and the average errors of the remaining reconstruction points are calculated to form a graph. If the uncertainty estimates align perfectly with the actual errors, then the reconstruction points with the highest errors are always removed first, which leads to the steepest decrease of the remaining error as the most uncertain points are being removed. This ideal sparsification curve is referred to as Oracle Sparsification. In practice, the uncertainty estimates do not align with the actual errors perfectly. Specifically, when a reconstruction point has a higher uncertainty but lower error compared to another point, the uncertainty estimates are considered not perfect. The area between the sparsification and its oracle indicates how different the uncertainty estimates are from the ideal ones, and can then be used to compute the Area Under Sparsification Error (AUSE). A smaller difference between the sparsification and its oracle results in a lower AUSE, and indicates that the uncertainty estimates are better because they align better with the actual errors.
@@ VI-B2 Results
When merging NeRF submaps, uncertainty-based filtering is particularly important. As shown in Fig. 9, the submaps can contain reconstruction artefacts which interfere with neighbouring submaps due to limited observation at the submap boundary. These artefacts can be however identified as they tend to have high epistemic uncertainty, and can be filtered to improve the merged reconstruction accuracy.
Figure 9: Comparison between two merged NeRF submaps with and without uncertainty filtering. On the left, the submaps contain artefacts which extend to neighbouring submaps and reduce overall reconstruction quality. These artefacts are mainly due to insufficient observations at the submap boundaries. On the right, we show how uncertainty filtering can be used to remove these artefacts which leads to a more accurate merged reconstruction.
Figure 10: Comparison between visual and lidar uncertainty. The reconstructions at the top are coloured by uncertainty estimates (red is high uncertainty, green is low uncertainty). Here, the lidar used was a Hesai XT32 which has a narrow FOV and cannot scan the upper part of the buildings, and the depth uncertainty estimates can identify such regions (red point clouds). This indicates that only visual information is used when reconstructing these areas.
The advantage of the disentangled uncertainty estimates is also demonstrated in Fig. 10 where we use a narrow vertical FOV lidar with wider vertical FOV cameras. Here, the lidar is not able to scan the upper part of the two buildings highlighted in Fig. 10, but the cameras can. Because of this, when we compute the epistemic uncertainty, we can see high lidar depth uncertainty for the upper part of both buildings. This indicates that the reconstruction is derived primarily from the visual data. In summary, from a sensor fusion point of view, our epistemic uncertainty estimation framework provides a systematic analysis of each sensor’s contribution to the final reconstruction.
@@ VI-C View Selection Strategy
In this section, we compare our visibility-based submapping strategy with an alternative distance-based submapping strategy proposed in [9]. As shown in Fig. 11, the building highlighted is divided into two submaps when using the distance-based submapping method. This is not ideal as it reduces the number of view constraints, which makes each submap’s partial reconstruction of that building have a lower quality. In addition, distance-based submapping puts the poses in A and C into the same submap, which is in fact not ideal. While pose A and pose C are spatially close, they have opposite viewing directions: pose A is looking at the highlighted building, while pose C is looking away from it. In comparison, visibility-based submapping moves pose A into the submap that contains the highlighted building, and pose C into another submap.
Figure 11: Comparison of two submapping strategies. Visibility information guides the clustering algorithm to group images looking at the same object together, which then leads to a more accurate and complete reconstruction of that object than algorithms that only consider distances.
@@ VI-D Multi-Camera Setup Ablation Study
The advantage of our multi-camera sensor setup is demonstrated qualitatively in Fig. 12. Compared to the three-camera setup, using only the front-facing camera leads to a reconstruction that is not only incomplete, but also with poorer geometry. Visual reconstruction with the photometric loss is limited to generating good-quality rendering only at the input viewing angle. The reconstruction using the front-only camera in Fig. 12 is trained with images looking in a single direction in the scene. This results in a poor geometric reconstruction when rendered from an unseen angle. In comparison, reconstruction with three cameras generates a more complete and more accurate reconstruction.
Figure 12: Comparison of reconstruction of HBAC building using the front camera only vs. using all three cameras. The three-camera setup generates more complete and accurate reconstructions compared to using only a single front-facing camera. The multi-camera setting is important in robotic applications where it would be infeasible to actively scan the entire scene to obtain strong viewpoint constraints.
@@ VI-E Effect of Bootstrapping SLAM Poses
The results are summarised in Tab. IV. For all COLMAP configurations, providing the SLAM prior poses not only accelerates pose computation, but also leads to better test rendering, compared to COLMAP without any initialisation. Our SLAM prior poses can also register all the images in the trajectory; meanwhile, COLMAP on its own only registers only 55%-95% images. Extracting more visual features per image (from 1024 to 8192) leads to a higher percentage of image registration and better visual reconstruction (PSNR and SSIM). This comes at the expense of a higher computation time, especially with the VocabTree matcher. Using the COLMAP Sequential Matcher is generally faster than Vocabulary Tree Matcher.
@@ VII Conclusions
Acknowledgements
The authors would like to thank Matias Mattamala for discussion, proofreading and helping with figures.